solve for x in simplest a+bi from x^2+8x+25=0

Answer by scott8148(6628)

completing the square ___ x^2 + 8x + 16 + 9 = 0 ___ (x + 4)^2 = -9

taking square root ___ x + 4 = ± 3i

subtracting 4 ___ x = -4 ± 3i